OpenStudy (anonymous):

Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. odd natural numbers B. rational numbers C. {0, 1} D. {0, 1, 2}

OpenStudy (vianne):

D. 2 -1 = 1 and 1 is in the {} 2 - 0 = 0, 0 is in the {} 1 -1 = 0 which is in {} C is also correct I think B is correct, I'm not 100% sure for that..

OpenStudy (anonymous):

Thank you Vianne I've been needing help 4 like 2 hours nobody was nice enough except you..... Thank you so very much

OpenStudy (zarkon):

1-2=-1

OpenStudy (vianne):

No Problem :)

OpenStudy (zarkon):

odd-odd =even example 5-3=2

OpenStudy (zarkon):

0-1=-1

OpenStudy (zarkon):

so A,C,D are not closed under subtraction